The nth term of a sequence is Tn=2+(n-2). Evaluate T3-T5
Answer by ikleyn(52793)

.

First simplify  Tn = 2 + (n-2) = 2 + n - 2 = n.


Now  T3 = 3, T5 = 5.


You do the rest.